From GlueXWiki
Jump to: navigation, search

FDC Weekly Meeting

Date: July 9, 2009

Participants: Daniel, Rich, Mark S., Simon, Bill, Beni, Mark I., Eugene

Next Meeting: July 16, 2009 @ 1:30 p.m.

General Construction Updates

 - The parts for the assembly jig on which the detector stack will be
   constructed are in our possession. Mark S. and Brian will assemble the
   jig when they have time and do some test fits.
 - Bill will take on the preparation for a stack assembly document. A
   draft will be prepared before assembly of the full-scale prototype and
   then fleshed out during the process with relevant photographs taken.
 - Cable preparations:
    * The LV cables have arrived and will be tested in the near future.
    * Fernando reported that the HV cables have been made and tested.
      He recommends that we test them with the new CAEN HV modules as
      there have been some recently reported problems with this type of
      module here at JLab.
    * The signal cables and connectors are in hand and Mark and Brian will
      begin construction as they have time.
    * We still need to decide how much of the outer jacket and braid will 
      be stripped off the signal cables.
 - Brian and Mark S. will begin making up the composite frames for the two
   gas windows. They should be ready within a couple of weeks.
 - Mark S. prepared the PO for the cables that will be used to attach to 
   the cathodes for the external ground. He will also order the sealing 
   epoxy that is used to cover the conducting epoxy.
 - We discussed some proposed modifications to the cathode tensioning 
   facility that Bill would like. As it turns out that the most recent
   cathode that we made with weights turned out quite well (i.e. no air
   bubbles). In light of this, we decided not to pursue this discussion
   too far and to revisit it after we have made a couple of more cathodes.
 - The parts for the second gusset ring have been delivered.

Wire Frame Update

 - Repair work on the second frame (P2) has been on hold as we lost our
   microscope and it has taken us some time to scramble a bit and find a
   suitable alternative. We are very close and Brian and Mark S. will get
   back on making repairs. Rich has started to look at the frame P4, which 
   was alledgedly the worst of the frames mechanically, and he claims that 
   it seems to have less problems than the others. Still awaiting full 
 - We have found that the boards are quite dirty and have quite a bit of
   solder flux. We are starting to develop cleaning plans and will ensure
   the boards are clean before stack assembly.
 - We need to have a dedicated meeting to design what the schedule for
   work on the Phase 3 wire winding will be. The highest priority for
   work scheduling will be doing work on upgrades of the wire winding
   facility. This includes a better strongback design, the design and
   construction of precision combs, modification to the wire winding
   table, and a list of other items with the system. Stay tuned, more
   to come.
 - We are pushing hard to begin wire frame construction early next year and
   begin wire winding before the end of 2010. Hopefully this can all procede.
   We are now working to complete the simulation studies with Simon to
   determine if we want to offset neighboring wire layers. More to come in
   the upcoming months.
 - The wire frame construction document is located at:
   /u/group/halld/Engineering/PRELIMINARY DOCUMENTS. We will include
   the information on the wire electroplating process shortly.
 - Bill is now designing the jig for the wire electroplating operation
   and it will be ready for submission to the shop at some point in the
   future. We will make a decision shortly if/when we will electroplate
   the wires for the full-scale prototype.
 - Brian received the SEM scans for our wire from the W&M folks. He will
   generate a final QA report with photos and test results for all wires
   in the next several weeks.

Cathode Update

 - Progress updates:
     * Mark S. has completed edge sealing and gas hole inclusion and sealing
       in 5 of the 8 cathode frames. He is working on this as he has time.
     * Cathode construction progress:
        - All 7 cathode panels (2 5-um Cu and 5 2-um Cu) have been prepared.
        - 3 cathodes (1 5-um Cu and 2 2-um Cu) have been tensioned and
          laminated to their laminate rings.
        - 1 cathode (5-um Cu) is in the tensioning facility now and has
          been tensioned. Presently studying the tension relaxation with
          time before completing the lamination process.
        - 1 cathode (2-um Cu) has been attached to the second transfer
          ring and is awaiting tensioning and lamination early next week.
        - 2 completed cathode panels are hanging in the clean tent.
     * Each cathode constructed this far has some sort of problem or
       non-ideal feature. Mark S. is keeping track of these in our log.
       One of our 7 cathode panels (a 2-um Cu) has a significant amount of
       epoxy +/-5 inches about each panel seam and most likely will not be
       usable. We are looking into way to try to remove the epoxy, but it
       will take some chemistry. We shall see.
     * We have a plan in place to get the completed cathode boards back under
       environmental control. Hopefully this will happen next week (but of
       course I said this last week and the week before and the week before
     * We have a bit of a back up doing the flatness scans of the completed
       cathodes. DSC will talk to Mark/David/Micah to find out where we
     * Mark S. tensioned the second 5-um cathode board this week, but he
       saw a 15-18% decrease the next day. We discussed how to proceed.
       The nominal tension on each board should be set to 500 N/m. We will
       allow the board to sit overnight and remeasure the tension the next
       day. If is is stable to within 10%, we will complete the lamination.
       If it is outside of this bound, we will retension and relaminate.
       Note that relative humidity should be recorded for each measurement.
 - The Fast Electronics group will complete the soldering of the rigid-flex
   assemblies within about 2.5 weeks. A small number will be made available
   on a shorter time frame for Mark S. to practice soldering them to the
   cathode with the epoxy on the front. Mark S. will be responsible (for now)
   for soldering the rigid-flex assemblies to the cathodes.
 - The current draft of the cathode construction document being worked on
   by Bill is located at /u/group/halld/Engineering/PRELIMINARY DOCUMENTS. 
   As new pictures are taken, they are being incorporated. As procedures 
   are modified or fine tuned, the document is being updated.
 - We are standing by waiting for information on the PR for the 2-um board
 - Roger needs to prepare a document for QA/stuffing/cleaning for the
   cathode boards and a similar document for the cathode daughter
   boards and ground boards.

Cooling System Tests

 - Fernando is in the process of writing up the results from his cooling 
   system studies.  Stay tuned for the GlueX note. He indicated that he
   is back to work on this document and it should be done shortly.

Cosmic Ray Test Stand

 - Beni is continuing work to setup the cosmic ray telescope and test
   system in EEL 126. Progress has been slow but steady with getting CODA
   and the DAQ working. There is still much to do, but things are coming
   along. In the past week, Beni has gotten a VXS VME crate from the Fast
   Electronics group and a 250-MHz JLab FADC. 
 - We still have some mysterious behavior with the small-scale prototype
   response with wiggles in the X vs. Y distributions that need to be
   understood. We decided to have a separate meeting to discuss what work
   we would like to do on the small-scale prototype.
 - Beni has started a bench test plan for the full-scale prototype and
   is presently working on it after incorporating feedback. It was also
   suggested to send it to Curtis for feedback. He will post it as a 
   GlueX-note shortly.
 - The machined parts for the support/rotation of the full-scale prototype
   in the cosmic ray stand have been delivered. We are awaiting some purchased
   parts to arrive. After we have everything in hand, it will be assembled
   and test fitted.
 - Safety paperwork for the cosmic ray test stand has been turned in for
   signatures, but the process is slow. Upon approval we will fill the 
   chambers with gas and be able to turn on the high voltages. At present,
   we have decided to locate the gas bottles outside of the EEL building.

Work List

 - The FDC short-term work list has been posted on the FDC web site.  
   This is continually being updated and DSC welcomes any feedback or 
   comments from the group.

Minutes prepared by Daniel. Send any comments or corrections along.